OpenStudy (anonymous):

Find an equation for the tangent line to the curve at point defined by the given value of t. x=3t^2+8, y=t^6, t=-2

OpenStudy (anonymous):

You differentiate both x and y with respect to t, then plug in t=-2

OpenStudy (anonymous):

i did that but somehow i am not getting the right answer

OpenStudy (anonymous):

what did you get as the derivatives?

OpenStudy (anonymous):

6t^5/6t=t^5/t

OpenStudy (anonymous):

well do X first, alone

OpenStudy (anonymous):

ok, it will be 6t

OpenStudy (anonymous):

right so substitute t=-2

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

now do the same for y

OpenStudy (anonymous):

6t^5=6(-2)^5=6*-32=-192

OpenStudy (anonymous):

so i get 16

OpenStudy (anonymous):

hmm, so, what does your book have as the answer?

OpenStudy (anonymous):

i have to write equation for the tangent line. So, i found my slope which is 16. For x and y do i evaluate the given equations at the value of t?

OpenStudy (anonymous):

yes, now that you have the slope you only need the intercept

OpenStudy (anonymous):

so if you put t=-2 into the original equations you get the point the curve passes through at t=-2. Now you have both a point and a slope, and that's enough to write the equation for a line.

OpenStudy (anonymous):

ok thanx ktklown.:)

OpenStudy (anonymous):

np :)